Need help, please show work for solutions.

1.) An investor just invested $10,000 in an investment that is expected to earn a 6% interest rate. Assuming the 6% annual return is realized, what will be the value of the investment at the end of 25 years?

2.) If you deposit $45,000 into a 5-year CD today earning 4% interest compounded quarterly, what would be the account balance be at the end of 5 years?

3.) A 22-year old college student has been promised a $1 million check at this 50th birthday (28 years from today). What is the present value of the $1 million today assuming an interest rate of 5%?

4.) An investor has been offered an investment opportunity that will pay him $25,000 every year for 10 years, with the first payment coming one year from today. What is the present value of the 10-year annuity at an annual interest rate of 8%?

5.) Jennifer has been given money for her birthday. She puts her money in a savings account offers an annual interest rate of 4.50%, compounded monthly. What is the effective annual rate (EAR) on the saving account?

6.) An individual invests $10,000 today in an investment that is expected to be worth $20,000 in 6 years. What annualized rate of return is the investor expecting to receive on the investment?

7.) A college graduate just bought a new house. The home was purchased with a 30-year loan of $350,000 with monthly payments at an interest rate of 4%. What is the amount of the monthly loan payment?

8.) An investor is evaluating an investment that pays $50,000 for 3 consecutive years and then $25,000 in year 4. The first $50,000 payment comes one year from today. What is the present value of the investment’s cash flows at an interest rate of 10%?
Answer #1

Answer 1:

PV = $10,000

Interest rate = 6%

Number of periods = 25

FV = FV (rate, nper, pmt, pv, type) = FV (6%, 25, 0, -10000, 0) = $42918.71

Value of the investment at the end of 25 years = $42,918.71

Answer 2:

PV = $45,000

Quarterly Interest rate = 4% / 4 = 1%

Number of periods = 5 * 4 = 20

FV = FV (rate, nper, pmt, pv, type) = FV (1%, 20, 0, -45000, 0) = $54908.55

Account balance be at the end of 5 years = $54,908.55

Answer 3:

FV = $1 million = $1,000,000

Annual Interest rate = 5%

Number of periods = 28

PV = PV (rate, nper, pmt, fv, type) = PV (5%, 28, 0, -1000000, 0) = $255093.64

Present value of the $1 million today = $255,093.64

Answer 4:

It is an ordinary annuity

PMT = $25,000

Annual Interest rate = 8%

Number of periods = 10

PV = PV (rate, nper, pmt, fv, type) = PV (8%, 10, -25000, 0, 0) = $167,752.03

Present value of the 10-year annuity = $167,752.03

As HOMEWORKLIB's policy, I have answered 4 questions